A ball bearing of radius 20cm and mass of 0.2kg moving through a liquid of viscosity 1.2×10^-6 NSM^-1. What is the terminal velocity?
1
Expert's answer
2020-11-16T07:36:48-0500

Given: Radius of bearing "r=20cm,=0.2m"

mass of bearing "m=0.2kg"


Viscosity of liquid "\eta=1.2\times10^{-6} Nsm^{-1}"


To find: Terminal velocity =?


Solution: we know that,

Mass =volume "\times" density

"0.2=\dfrac{4}{3}\pi(0.2)^3\times" Density


"\Rightarrow" Density"\rho=\dfrac{3}{4\pi\times 0.04}" =5.965kg/"m^3"


Terminal velocity is given by-

"v_t=\dfrac{2gr^2}{9\eta}\rho"


"\Rightarrow v_t=\dfrac{2\times 9.8\times (0.2)^2\times 5.965}{9\times 1.2\times 10^{-6}}"


"\Rightarrow v_t=\dfrac{4.646\times 10^6}{10.8}=0.43\times 10^6ms^{-1}"


Hence the terminal velocity of object is "0.43\times 10^6ms^{-1}"